OpenStudy (anonymous):

Solve the system using the addition method. 5x - 2y = 3 2x - y = 0

OpenStudy (cwrw238):

multiply second equation by -2 adding will then give an equation in x only

OpenStudy (cwrw238):

solve this and plug in value of x into one of the original equations to find value of y

OpenStudy (anonymous):

how do i solve the second by 2

OpenStudy (anonymous):

-4x+2=0 is that right?

OpenStudy (anonymous):

the answers 3?

OpenStudy (cwrw238):

nearly its -4x + 2y = 0 now adding this to 5x - 2y = 3 will eliminate y

OpenStudy (anonymous):

1x+0y=3

OpenStudy (cwrw238):

5x - 2y = 3 -4x + 2y = 0 add x = 3 now plug x=3 into 2x - y = 0 2(3) - y = 0 6 = y x=3,y=6

OpenStudy (anonymous):

so the y value would be 6

OpenStudy (anonymous):

Thanks! also i need help on this one pleaseeee To solve this system using the addition method, you would need to multiply the first equation by what number in order for the y's to add out? 3x - y = 3 -2x + 2y = 6